# fev-bench: A Realistic Benchmark for Time Series Forecasting

### v1

@@ D.2 Pairwise comparison
Figure 4: Pairwise win rates $W_{jk}$ (Equation 8) of all models against each other under the scaled quantile loss (SQL) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better.
Figure 5: Pairwise skill scores $S_{jk}$ (Equation 9) of all models against each other under the scaled quantile loss (SQL) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Note that pairwise skill score is not symmetric, $S_{jk}\neq S_{kj}$.
Figure 6: Pairwise win rates $W_{jk}$ (Equation 8) of all models against each other under the mean absolute scaled error (MASE) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better.
Figure 7: Pairwise skill scores $S_{jk}$ (Equation 9) of all models against each other under the mean absolute scaled error (MASE) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Note that pairwise skill score is not symmetric, $S_{jk}\neq S_{kj}$.
@@ E.2 Evaluation results
Model Avg. win rate (%) Skill score (%) Median runtime (s) Leakage (%) # failures
TiRex 84.6 43.4 1.3 0 0
TimesFM-2.5 79.0 44.0 76.1 5 0
Toto-1.0 77.9 43.0 66.3 5 0
TabPFN-TS 71.2 46.2 275.0 0 0
Moirai-2.0 69.6 41.7 1.6 30 0
Chronos-Bolt 66.2 40.4 1.1 0 0
Sundial-Base 53.1 37.6 23.1 0 0
Stat. Ensemble 48.8 27.3 726.8 0 2
AutoARIMA 47.3 30.6 226.0 0 2
AutoETS 35.0 -9.4 14.8 0 0
AutoTheta 27.7 8.5 8.1 0 0
Seasonal Naive 17.7 0.0 2.3 0 0
Naive 13.8 -36.9 2.2 0 0
Drift 8.1 -36.0 2.2 0 0
@@
Model Avg. win rate (%) Skill score (%) Median runtime (s) Leakage (%) # failures
TiRex 78.8 31.4 1.3 0 0
TimesFM-2.5 77.9 32.3 76.1 5 0
Toto-1.0 75.6 30.9 66.3 5 0
TabPFN-TS 66.9 35.1 275.0 0 0
Moirai-2.0 66.2 30.1 1.6 30 0
Chronos-Bolt 64.2 28.4 1.1 0 0
Sundial-Base 61.9 29.8 23.1 0 0
Stat. Ensemble 50.8 21.1 726.8 0 2
AutoARIMA 44.2 20.4 226.0 0 2
AutoTheta 35.4 14.8 8.1 0 0
AutoETS 33.8 2.3 14.8 0 0
Seasonal Naive 16.2 0.0 2.3 0 0
Naive 15.0 -17.4 2.2 0 0
Drift 13.1 -17.2 2.2 0 0

### v2

@@ D.2 Pairwise comparison
Figure 4: Pairwise win rates $W_{jk}$ (Eq. 6) of all models against each other under the scaled quantile loss (SQL) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better.
Figure 5: Pairwise skill scores $S_{jk}$ (Eq. 7) of all models against each other under the scaled quantile loss (SQL) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Note that pairwise skill score is not symmetric, $S_{jk}\neq S_{kj}$.
Figure 6: Pairwise win rates $W_{jk}$ (Eq. 6) of all models against each other under the mean absolute scaled error (MASE) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better.
Figure 7: Pairwise skill scores $S_{jk}$ (Eq. 7) of all models against each other under the mean absolute scaled error (MASE) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Note that pairwise skill score is not symmetric, $S_{jk}\neq S_{kj}$.
@@ E.2 Evaluation results
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 91.8 51.1 1.2 0 0
TiRex 80.4 43.4 0.4 0 0
TimesFM-2.5 74.8 44.0 5.2 5 0
Toto-1.0 74.1 43.0 44.3 5 0
TabPFN-TS 67.1 46.2 176.3 0 0
Moirai-2.0 65.0 41.7 0.4 30 0
Chronos-Bolt 61.4 40.4 0.5 0 0
Sundial-Base 50.0 37.6 8.3 0 0
Stat. Ensemble 45.7 27.3 261.8 0 2
AutoARIMA 44.3 30.6 79.6 0 2
AutoETS 32.9 -9.4 6.4 0 0
AutoTheta 25.7 8.5 5.1 0 0
Seasonal Naive 16.4 0.0 0.8 0 0
Naive 12.9 -36.9 0.8 0 0
Drift 7.5 -36.0 0.8 0 0
@@
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 88.9 40.4 1.2 0 0
TiRex 75.0 31.4 0.4 0 0
TimesFM-2.5 73.7 32.3 5.2 5 0
Toto-1.0 72.3 30.9 44.3 5 0
TabPFN-TS 63.6 35.1 176.3 0 0
Moirai-2.0 62.1 30.1 0.4 30 0
Chronos-Bolt 59.6 28.4 0.5 0 0
Sundial-Base 58.6 29.8 8.3 0 0
Stat. Ensemble 48.2 21.1 261.8 0 2
AutoARIMA 41.4 20.4 79.6 0 2
AutoTheta 33.2 14.8 5.1 0 0
AutoETS 31.8 2.3 6.4 0 0
Seasonal Naive 15.0 0.0 0.8 0 0
Naive 13.9 -17.4 0.8 0 0
Drift 12.5 -17.2 0.8 0 0